Find all solutions of the given differential equations and then find the particular solutions for which a point (x,y) is given:


dy/dx = x


1
Expert's answer
2021-09-05T14:23:00-0400

First we solve the integral by separating the terms:


"\frac{dy}{dx}={x} \to \intop dy=y=\int{xdx}=\frac{x^{2}}{2}+C
\\\implies y=\frac{x^{2}}{2}+C"


Then, we substitute the coordinates (xi,yi) to find C for the particular solution:


"y_i=\frac{x_i^{2}}{2}+C \implies C=y_i-\frac{x_i^{2}}{2}"


In conclusion:


"\\text{General solution: } \\\\ y=\\frac{x^{2}}{2}+C \\\\ \\text{Particular solution } (x_i,y_i):\\\\ y=\\frac{x^{2}}{2}+y_i-\\cfrac{x_i^{2}}{2}\n\n\u200b"